OpenStudy (anonymous):

Write the equation of direct variation for x = -5 and y = 90

OpenStudy (anonymous):

It's easier than you may seem. Recall the direction variation equation I gave you earlier: \[\frac{ y }{ x }=k\] Just plug in your numbers and you have your equation. Some teachers can be picky about which form they'd like it in, so I would write it in both \[\frac{ y }{ x }=k\] AND \[y=kx\]

OpenStudy (anonymous):

okay i got -18

OpenStudy (anonymous):

is is 90=(-18)-5

OpenStudy (anonymous):

Since it asks just for the equation, you don't even half to solve. Just plug in the numbers into the equation and that's it.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

so just put it like this 90=k-5

OpenStudy (anonymous):

If you mean multiply k by (-5) and not subtract, then yes.

OpenStudy (anonymous):

kk XD

OpenStudy (anonymous):

But some teachers can get picky about which form they want, so I'd plug it into both forms.